the population standard deviation is estimated by the sample standard deviation of eight hours what is the probability that the difference between the sample mean and the population mean


1
Expert's answer
2021-06-01T08:42:32-0400

For example:

The estimated difference between the sample mean and the population mean:

"\Delta=3"

Sample:

"n=100"

The sample standard deviation:

"\sigma=8"

The sample mean:

"\overline{x}=20"

The population mean:

"\mu=15"

Significance level:

"\alpha=5\ \%"


"H_0:" "\Delta>3"

"H_a\le3"


Then:


"z=\frac{\overline{x}-\mu-\Delta}{\sigma/\sqrt{n}}=\frac{20-15-3}{8/\sqrt{100}}=0.03"


p-value:

"P(\Delta>3)=1-P(z<0.03)=1-0.51197=0.48803"

Since p-value > "\alpha" , we accept the null hypothesis.